OpenStudy (anonymous):

slope intercept form of (3,4),(1,-2)

OpenStudy (anonymous):

y=3x-5 do u need an explanation

OpenStudy (anonymous):

no haha

OpenStudy (anonymous):

could you help me with another one.?!

OpenStudy (anonymous):

yeah sure

OpenStudy (anonymous):

still need slope intercept form: slope = -1/2 and the ordered pair is (0,-9)

OpenStudy (anonymous):

y=-1/2x-9 explanation?

OpenStudy (anonymous):

no! but another?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

slope = 1 ordered pair is (-1,-4)

OpenStudy (anonymous):

y=x-3

OpenStudy (anonymous):

thank you so muchh!!

OpenStudy (anonymous):

ur welcome